# Hyperfine spectroscopy of optical-cycling transitions in singly ionized thulium

### v1

@@ III Results
The $A$ constants for the ground state and the $236.95$ and 31 926.8231\,926.82-cm-1 states were obtained from collinear laser spectroscopy (CLS) measurements in Darmstadt [Bodnar.2023] and were used as starting values for determining the remaining constants. The HFS $A$ constants of the two low lying states from CLS as well as the $A$ constant of the $22308.82$-cm-1 state, determined in this work, agree within $1\sigma$ with the values available in the literature [Kebapci.2024].
IV Proposed cooling schemes
This section presents the most promising laser cooling schemes, selected based on their transition strengths and the accessibility of the required wavelengths with current laser technology.
IV.1 Laser cooling at 450 nm
Figure 4: (a) Solution of the rate equations for the 450-nm cycle. State labels are given in cm-1. Dashed lines are without repumping lasers, solid lines include $749$, $846$ and $855$-nm repumping lasers. Unknown Einstein $A_{21}$-coefficients were taken from the DREAM database [Quinet.1999, Quinet.DREAM, Palmeri.2025]. (b) Experimental scattering rate without the infrared repumping lasers at a He pressure of $4.7\cdot 10^{-6}$ Torr. Blocking the $450$-nm lasers for 22 minutes pauses the optical population transfer. (c) Increase of the steady-state scattering rate with He-buffer-gas pressure.
When searching for a closed optical cycle through the ground state at a convenient wavelength, transitions near $450$ nm appear to be an attractive choice. Transitions at $448$ and $453$ nm connect the even-parity 22 308.8222\,308.82-cm-1 state ($4f^{12}5d6s$, $J=4$) to the odd-parity ground and $236.95$-cm-1 states, respectively. Results in the DREAM database suggest that the 22 308.8222\,308.82-cm-1 state could also decay via an $E1$ transition at $749$ nm to an odd-parity state at 8 957.478\,957.47-cm-1[Quinet.1999, Quinet.DREAM, Palmeri.2025]. Because this state has the same parity as the ground states, $E1$ decays back to the cooling cycle are forbidden. Based on the electron $g$ factor and angular momentum geometry, we estimate a lifetime of $\approx\!90$ ms for this state, which implies the need for repumping lasers. In practice, however, we did not observe any evidence of population accumulating in this state. In contrast, we do observe a slow population transfer, on the order of seconds, to a long-lived $4f^{12}6s^{2}\,{}^{3}\mathrm{H}_{6}$ state at 12 457.2912\,457.29 cm-1, which can be repumped with $846.31$, $1152.58$, or $1286.75$-nm light. These repumping wavelengths correspond to transitions to the 24 273.2024\,273.20 $(\nicefrac{{7}}{{2}},\nicefrac{{5}}{{2}})_{5}^{\mathrm{o}}$], 21 133.6821\,133.68 $(\nicefrac{{7}}{{2}},\nicefrac{{5}}{{2}})_{6}^{\mathrm{o}}$], and 20 228.7520\,228.75-cm-1 $(\nicefrac{{7}}{{2}},\nicefrac{{3}}{{2}})^{\mathrm{o}}_{5}$] states, respectively. Based on the observed HFS spectra, the $846$-nm and $1153$-nm transitions appear to be stronger repumping channels than the $1287$-nm transition.
[…]
We find that adding helium buffer gas greatly increases the observed signal in both 313 and 450-nm cycles. This experimental observation is consistent with the cycles being open, as the buffer gas could be quenching undiscovered, long-lived leakage channels. A second possibility is that the buffer gas simply aids in trapping, leading to a higher ion density. The problem of leakage in the 450-nm cycle is further illustrated by a rate-equation simulation performed with the qspec Python package [qspec]. As is shown in FIG. 4(a), it predicts rapid population loss without repumping. Without the infrared repumping lasers, the population is first driven into the $8957.47$-cm-1 and 20 228.7520\,228.75-cm-1 states on a sub-millisecond timescale. Both states continuously decay back into the ground state, while about $38$ % of the 20 228.7520\,228.75-cm-1 state branches into the long-lived $12457.29$-cm-1 state on a timescale of $20$ s. The $22308.82$-cm-1 state, which is the only state contributing to the fluorescence signal, is not populated significantly after approximately $5$ ms. However, a small amount of less than $1$ % remains until the $12457.29$-cm-1 state is populated. These simulations qualitatively agree with the experimental findings in FIG. 4(b), showing the slowly decaying fluorescence signal from this remaining population when the pump lasers are on, but no population loss when the light is blocked. The 20 228.7520\,228.75-cm-1 state being the bottleneck explains why no effect of the $749$-nm laser on the fluorescence signal is observed when only the $12457.29$-cm-1 state is repumped. Consequently, the full optical cylcling scheme is expected to require $749$, $846$ and $855$-nm repumping lasers, reaching steady-state population after about $30$ ms, with $23$ % of the population contributing to the fluorescence signal. Discrepancies between simulation and experiment in the population transfer timescale and observed scattering rate suggest that additional $M1$ decay channels, not listed in the DREAM database, might need to be considered. FIG. 4(c) shows the increased fluorescence signal when using He-buffer gas, here measured with the $313$-nm transition; hinting at quenching of the metastable states.
IV.2 Laser cooling at 313 nm
The $313$-nm line is the strongest transition we measured, with an Einstein coefficient of $A_{21}=2\pi\times 16.9$ MHz [NIST]. It connects the ground state to an even-parity $J=5$ state with a $4f^{12}5d6s$ electronic configuration at 31 926.8231\,926.82 cm-1. Its relatively large linewidth, together with its short transition wavelength, offers a high Doppler cooling rate.
Based on $E1$ selection rules, some population from the $313$-nm cooling cycle is expected to decay into the 22 457.5122\,457.51-cm-1 $(\nicefrac{{7}}{{2}},\nicefrac{{3}}{{2}})_{4}^{\mathrm{o}}$] and 23 524.0923\,524.09-cm-1 $(\nicefrac{{7}}{{2}},\nicefrac{{5}}{{2}})_{4}^{\mathrm{o}}$] levels. We applied light at $854.69$, $904.43$, $1056.04$, and $1190.09$ nm to repump population through different excited states within the $313$-nm cycle (see FIG. 2), but observed no measurable change in fluorescence. This indicates that either the corresponding decay channels are weaker than the $854.84$-nm channel, or the lifetimes of these intermediate states are too short to cause significant population trapping.
Another potential metastable $(6,1)_{5}$ state suggested by the DREAM database [Quinet.1999, Quinet.DREAM, Palmeri.2025] is located at 16 567.4716\,567.47 cm-1. Based on its energy and selection rules, this level should decay to the ground state via an EE1 transition at approximately $604$ nm. However, driving the 1 297.741\,297.74-nm transition connecting this state to the 24 273.2024\,273.20-cm-1 level produced no observable change in the fluorescence signal.
IV.3 Laser cooling on other transitions
A $370$-nm transition connecting the ground state to a $J=4$ state at 27 254.4227\,254.42 cm-1 (not shown in FIG. 2) is another candidate for laser cooling. While this was the first transition we observed in the HCL, it is estimated to be $5$ times weaker than the $313$-nm transition. Furthermore, it shares most of the decay pathways to higher-$J$ states and introduces new decay channels toward $J\leq 3$ metastable states. The $313$-nm line is therefore considered a more feasible option.
While this work has focused on transitions from the ground state to $J=4$ and $J=5$ manifolds, other configurations are possible. For example, a recent UV spectroscopy study [DenHartog.2024] identified several promising transitions originating from the 8 957.478\,957.47-cm-1 state, including transitions to $J=3$ states near $363$ and $331$ nm with high branching factors of $0.995$ and $0.971$, respectively. Although lasers at these specific wavelengths were not available for the present work, these transitions represent a compelling avenue for future investigation.
IV.4 Sympathetic cooling with $^{138}\text{Ba}^{+}$
Our ion trap can reliably trap and laser cool large clouds of $^{138}\text{Ba}^{+}$ ions. By repeatedly switching the trap potential between high and low values, ions can be selectively removed to produce a laser-cooled few-ion crystal or a single ion. In principle, simultaneously co-trapping $\text{Ba}^{+}$ and $\text{Tm}^{+}$ should allow for the sympathetic cooling of $\text{Tm}^{+}$ ions during $\text{Ba}^{+}$ laser cooling. While we have demonstrated co-trapping of large $\text{Ba}^{+}$ and $\text{Tm}^{+}$ clouds via trap-door loading, attempts to crystallize the two species simultaneously, so far, resulted in the loss of all $\text{Tm}^{+}$ ions. An alternative approach is to first trap and crystallize $\text{Ba}^{+}$ and then introduce $\text{Tm}^{+}$ without lowering the trap potential, either by resonant photo-ionization of neutral Tm or by capturing $\text{Tm}^{+}$ through collisions with the trapped $\text{Ba}^{+}$ ions.
[…]
V 12457.29 level as a qubit candidate
Figure 5: The decay curve, pump-and-probe experiment, microwave spectrum and magnetic substructure of the $12457.29$-cm-1 $^{3}\mathrm{H}_{6}$] state. (a) Population in this state as a function of time. An exponential function was fitted to the data, yielding a decay constant of $5.2(3)$ min into the $^{3}\mathrm{F}_{4}$ ground state through $M2$ or $E3$ decay. (b) Example of a pump-and-probe measurement used to determine the lifetime of the $12457.29$-cm-1 state. (c) Zeeman-resolved microwave spectrum of the transitions between the two hyperfine states of the $12457.29$-cm-1 level, measured at a magnetic field of $1.34$ G. A sum-of-Gaussians model was fitted to the data, with peak widths proportional to the peak’s Zeeman shifts, to account for the variation of the magnetic field across the ion cloud. (d) Pairwise degenerate transitions, which contribute to the twelve inner peaks in the spectrum, are shown as solid lines, with the degenerate pair having the same color. Non-degenerate transitions, which contribute to the twelve outer peaks, are shown as dashed gray lines.
@@ IV.4 Sympathetic cooling with $^{138}\text{Ba}^{+}$
We observe experimentally that both of the optical cycles investigated lead to population leaking into the 12 457.2912\,457.29-cm-1 state, which makes this state a natural choice for the metastable state in the OMG protocol. From now on, we will refer to this state as the “Gollum” state. So far, the fastest way to transfer the ground-state population to this level is by applying $313$-nm light, which pumps the population into the state via its strong leakage channels. A population imbalance between the $F=\nicefrac{{11}}{{2}}$ and $\nicefrac{{13}}{{2}}$ sublevels can be created by coupling one of them to a different level, e.g., with $846$, $1153$, or $1287$-nm light. Addressing both HFS sublevels with these lasers will fully recover the population from this level.
@@
To test if the Gollum state lives long enough to be used as a qubit state, we measured its decay time using a pump-probe approach, with the result shown in FIG. 5(a). In this measurement, the $313$-nm laser is used to “pump” the population into the metastable state, while the closed $450$-nm cycle serves as the “probe.” The $846$-nm laser is used as a repumper to return the population from the Gollum state back into the $450$-nm probe cycle for detection.
The experimental sequence is shown in FIG. 5(b). At the beginning, $^{169}\text{Tm}^{+}$ is trapped through laser ablation. Applying the $313$-nm laser decreases the detected photon rate down to the background level, pumping ions out of the 450-nm cycle into the Gollum state. Then, all lasers are turned off for a varying dark time, during which the population spontaneously decays into the ground state. To measure the fraction of the population that remained in the metastable level, the $450$-nm lasers are first turned back on, probing the decayed population. After a few seconds, $846$-nm light is added to recover all the population to the $450$-nm cycle. The change in fluorescence upon adding $846$-nm light, normalized to the maximum fluorescence from all ions, yields the fractional population in the Gollum state. Note that for the normalization, it was taken into account that the same number of ions produced about 10 % more fluorescence in the $450$-nm cycle when the $846$-nm laser was added, and that there was a slow loss of ions. Fitting an exponential decay model to the resulting decay curve yields a minimum lifetime of $5.2(3)$ min. For the fit, it was assumed that immediately after turning off the $313$-nm laser, all of the population is in the Gollum state and for an infinite time after, all of the population has decayed back to the ground state. The determined lifetime is sufficiently long to prepare a qubit and execute quantum gates [DiVincenzo.2000, Blatt.2004].
@@
The high angular momentum of $J=6$ enables the implementation of absorption-emission codes (Æ codes), as proposed in [Jain.2024]. These codes are robust against first-order absorption and emission errors but require a long state lifetime for operation. For the preparation of any qubit, the magnetic sublevels of the state need to be considered, which we analyzed with microwave spectroscopy. For this, the $1152$-nm laser was resonant with either of the $\nicefrac{{11}}{{2}}\rightarrow\nicefrac{{11}}{{2}},\nicefrac{{13}}{{2}}$ transitions, such that only the $\nicefrac{{11}}{{2}}$ state is repumped through the 21 133.6821\,133.68 cm-1 level. The fluorescence signal is taken from the light scattered in the $450$-nm cycle, such that increasing the population recovery rate from the Gollum state increases the signal. By applying resonant microwaves, the previously unaddressed population in the $F_{+}=\nicefrac{{13}}{{2}}$ state is transferred into the $F_{-}=\nicefrac{{11}}{{2}}$ state, where it can be repumped by the $1152$-nm laser. The resulting spectrum is shown in FIG. 5(c) with the hyperfine structure splitting between the $F_{-}=\nicefrac{{11}}{{2}}$ and $F_{+}=\nicefrac{{13}}{{2}}$ state of Gollum being $2.55$ GHz, with $\nicefrac{{11}}{{2}}$ having a higher energy. The separable and strong microwave resonances show that the Gollum state is well-suited for microwave-based gates.
@@
where $g_{J}$ is the electronic $g$-factor. The fact that $\nicefrac{{g_{-}}}{{g_{+}}}=\nicefrac{{(J+1)}}{{J}}$ leads to peculiar degeneracies in the microwave transitions between the $m_{F}$ states of the two HFS levels: each $\pi$-transition $m^{\pi}\rightarrow m^{\pi}$ is degenerate with a single $\sigma^{\pm}$ transition $m^{\pm}\rightarrow m^{\pm}\pm 1$ for which $m^{\pi}=m^{\pm}\pm J$. On top of that, the degenerate transitions align by strength: the strongest $\pi$ transitions $m_{F}=\pm\nicefrac{{1}}{{2}}\rightarrow m_{F}^{\prime}=\pm\nicefrac{{1}}{{2}}$ are degenerate with the strongest $\sigma$ transitions $m_{F}=\mp\nicefrac{{13}}{{2}}\rightarrow m_{F}^{\prime}=\mp\nicefrac{{11}}{{2}}$; the second strongest $\pi$ transitions $m_{F}=\pm\nicefrac{{3}}{{2}}\rightarrow m_{F}^{\prime}=\pm\nicefrac{{3}}{{2}}$ are degenerate with the second strongest $\sigma$ transitions $m_{F}=\mp\nicefrac{{11}}{{2}}\rightarrow m_{F}^{\prime}=\mp\nicefrac{{9}}{{2}}$, etc, which is illustrated in FIG. 5(d). This results in $2\times(2F_{-}+1)=24$ distinct peaks, with the peaks corresponding to the weakest transitions barely visible. The lower observed heights of the central peaks corresponding to the strongest transitions still remain a puzzle.
@@ VI Future directions
To increase the laser-cooling rate, first, the already known leakage channels of the $313$-nm cycling scheme can be closed, using two frequency tones at $846$, $854.8$, $927/944$ and $1307$ nm simultaneously. Even if this turns out to be insufficient, it will facilitate the search for the remaining repumping transitions as the population will be diverted into the unknown weak decay channels. In the $450$-nm cycle, the $749$-nm transition is the only feasible unknown decay channel and it is expected to be strong enough to be relevant [Quinet.1999, Quinet.DREAM, Palmeri.2025]. While He-buffer-gas quenching helped finding unknown repumping transitions, in the case of the $749$-nm transition, the estimated lifetime of its lower state ($8957.47$ cm-1) of $\approx\!90$ ms might be reduced below the mean time of  $\approx\,0.63$ ms the upper state takes to decay into it. Consequently, adding the repumping laser in He-buffer gas would not improve the active optical cycling time, so that finding both hyperfine components simultaneously in a vacuum setup would be required.
VII Conclusions
In conclusion, we have mapped out the hyperfine structure (HFS) required to assemble operational laser-cooling cycles in $^{169}\text{Tm}^{+}$, combining absorption spectroscopy in a hollow-cathode lamp with fluorescence spectroscopy in an ion trap. The data include hyperfine $A$ constants for key excited states, validate a strong $313$-nm cycling transition, and identify a practical 448448/$453$-nm fluorescence-probing and optical-cycling scheme. These optical cycles are supported by the near-infrared repumping transitions identified in this work.

### v2

@@ Results
The $A$ constants for the ground state and the $236.95$ and 31 926.8231\,926.82-cm-1 states were obtained from collinear laser spectroscopy (CLS) measurements at Technische Universität Darmstadt [4] and were used as starting values for determining the remaining constants. The HFS $A$ constants of the two low lying states from CLS as well as the $A$ constant of the $22308.82$-cm-1 state, determined in this work, agree within $1\sigma$ with the values available in the literature [15]. All HFS splittings lie between 11-55 GHz, making $\text{Tm}^{+}$ an ideal candidate for microwave-based qubit operations with well-separated HFS states.
12457.29-cm-1 level as a qubit candidate
@@
We observe experimentally that both of the optical cycles investigated lead to population leaking into the 12 457.2912\,457.29-cm-1 state, which makes this state a natural choice for the metastable state in the omgomg qubit protocol. From now on, we will refer to this state as the “Gollum” state. So far, the fastest way to transfer the ground-state population to this level is by applying $313$-nm light, which pumps the population into the state via its strong leakage channels. A population imbalance between the $F=\nicefrac{{11}}{{2}}$ and $\nicefrac{{13}}{{2}}$ sublevels can be created by coupling one of them to a different level, e.g., with $846$, $1153$, or $1287$-nm light. Addressing both HFS sublevels with these lasers will fully recover the population from this level.
Figure 4: Microwave spectroscopy of the 12457.29-cm-1 state. The decay curve, pump-and-probe experiment, microwave spectrum and magnetic substructure of the $12457.29$-cm-1 $^{3}\mathrm{H}_{6}$] state (Gollum). (a) Population in this state as a function of time. An exponential function was fitted to the data, yielding a decay constant of $5.2(3)$ min into the $^{3}\mathrm{F}_{4}$ ground state through $M2$ or $E3$ decay. (b) Example of a pump-and-probe measurement used to determine the lifetime of the $12457.29$-cm-1 state. The numbers in parentheses refer to the sequence steps described in Methods. (c) Zeeman-resolved microwave spectrum of the transitions between the two hyperfine states of the $12457.29$-cm-1 level, measured at a magnetic field of $1.34$ G. A sum-of-Gaussians model was fitted to the data, with peak widths proportional to the peak’s Zeeman shifts, to account for the variation of the magnetic field across the ion cloud. (d) Pairwise degenerate transitions, which contribute to the twelve inner peaks in the spectrum, are shown as solid lines, with the degenerate pair having the same color and labeled by the same geometric features, e.g., $\nicefrac{{13}}{{2}}\rightarrow\nicefrac{{11}}{{2}}$ is degenerate with $\nicefrac{{1}}{{2}}\rightarrow-\nicefrac{{1}}{{2}}$. Non-degenerate transitions, which contribute to the twelve outer peaks, are shown as dashed gray lines.
@@
To test if the Gollum state lives long enough to be used as a qubit state, we measured its decay time using a pump-probe approach, with the result shown in Figure 4(a). In this measurement, the $313$-nm laser is used to “pump” the population into the metastable state, while the closed $450$-nm cycle serves as the “probe”. The $846$-nm laser is used as a repumper to return the population from the Gollum state back into the $450$-nm probe cycle for detection. The fractional population of the Gollum state at the time of the repumping process is equal to the fractional “recovered signal” shown on the $y$-axis of Figure 4(a) and given by Eq. (4) in Methods. This measurement scheme is shown in Figure 4(b) and described in more detail in Methods.
Fitting an exponential decay model to the resulting decay curve yields a minimum lifetime of $5.2(3)$ min. For the fit, it was assumed that immediately after turning off the $313$-nm laser, all of the population is in the Gollum state and for an infinite time after, all of the population has decayed back to the ground state. The determined lifetime is sufficiently long to prepare a qubit and execute quantum gates [7, 3].
@@
The high angular momentum of $J=6$ is a sufficient condition for the implementation of absorption-emission codes (Æ codes), as proposed in Refs. [13, 2]. In addition, the Gollum doublet contains an $F=\nicefrac{{13}}{{2}}$ state, which is the minimum required for a the Clifford-friendly spin code described in Ref. [10], which has been shown to be an absorption-emission code [2, 13]. These codes are robust against first-order absorption and emission errors but require a long state lifetime for operation. For the preparation of any qubit, the magnetic sublevels of the state need to be considered, which we analyzed with microwave spectroscopy. For this, the $1153$-nm laser was resonant with either of the $\nicefrac{{11}}{{2}}\rightarrow\nicefrac{{11}}{{2}},\nicefrac{{13}}{{2}}$ transitions, such that only the $\nicefrac{{11}}{{2}}$ state is repumped through the 21 133.6821\,133.68 cm-1 level. The fluorescence signal is taken from the light scattered in the $450$-nm cycle, such that increasing the population recovery rate from the Gollum state increases the signal. By applying resonant microwaves, the previously unaddressed population in the $F_{+}=\nicefrac{{13}}{{2}}$ state is transferred into the $F_{-}=\nicefrac{{11}}{{2}}$ state, where it can be repumped by the $1153$-nm laser. The resulting spectrum is shown in Figure 4(c) with the hyperfine structure splitting between the $F_{-}=\nicefrac{{11}}{{2}}$ and $F_{+}=\nicefrac{{13}}{{2}}$ state of Gollum being $2.55$ GHz and $\nicefrac{{11}}{{2}}$ having a higher energy. The separable and strong microwave resonances show that the Gollum state is well-suited for microwave-based gates.
@@
where $g_{J}$ is the electronic $g$-factor. The fact that $\nicefrac{{g_{-}}}{{g_{+}}}=\nicefrac{{(J+1)}}{{J}}$ leads to peculiar degeneracies in the microwave transitions between the $m_{F}$ states of the two HFS levels: each $\pi$-transition $m^{\pi}\rightarrow m^{\pi}$ is degenerate with a single $\sigma^{\pm}$ transition $m^{\pm}\rightarrow m^{\pm}\pm 1$ for which $m^{\pi}=m^{\pm}\pm J$. On top of that, the degenerate transitions align by strength: the strongest $\pi$ transitions $m_{F}=\pm\nicefrac{{1}}{{2}}\rightarrow m_{F}^{\prime}=\pm\nicefrac{{1}}{{2}}$ are degenerate with the strongest $\sigma$ transitions $m_{F}=\mp\nicefrac{{13}}{{2}}\rightarrow m_{F}^{\prime}=\mp\nicefrac{{11}}{{2}}$; the second strongest $\pi$ transitions $m_{F}=\pm\nicefrac{{3}}{{2}}\rightarrow m_{F}^{\prime}=\pm\nicefrac{{3}}{{2}}$ are degenerate with the second strongest $\sigma$ transitions $m_{F}=\mp\nicefrac{{11}}{{2}}\rightarrow m_{F}^{\prime}=\mp\nicefrac{{9}}{{2}}$, etc, which is illustrated in Figure 4(d). This results in $2\times(2F_{-}+1)=24$ distinct peaks, with the peaks corresponding to the weakest transitions barely visible. The lower observed heights of the central peaks corresponding to the strongest transitions still remain a puzzle.
@@ Discussion
To increase the laser-cooling rate, first, the already known leakage channels of the $313$-nm cycling scheme can be closed, using two frequency tones at $846$, $854.8$, $927/944$ and $1307$ nm simultaneously. Even if this turns out to be insufficient, it will facilitate the search for the remaining repumping transitions as the population will be diverted into the unknown weak decay channels. In the $450$-nm cycle, the $749$-nm transition is the only feasible unknown decay channel and it is expected to be strong enough to be relevant [32, 33, 28]. While He-buffer-gas quenching helped finding unknown repumping transitions, in the case of the $749$-nm transition, the estimated lifetime of its lower state ($8957.47$ cm-1) of $\approx\!90$ ms might be reduced below the mean time of  $\approx\,0.63$ ms the upper state takes to decay into it. Consequently, adding the repumping laser in He-buffer gas would not improve the active optical cycling time, so that finding both hyperfine components simultaneously in a vacuum setup would be required.
Conclusions
In conclusion, we have mapped out the hyperfine structure (HFS) required to assemble operational laser-cooling cycles in $^{169}\text{Tm}^{+}$, combining absorption spectroscopy in a hollow-cathode lamp with fluorescence spectroscopy in an ion trap. The data include hyperfine $A$ constants for key excited states, validate a strong $313$-nm cycling transition, and identify a practical 448448/$453$-nm fluorescence-probing and optical-cycling scheme. These optical cycles are supported by the near-infrared repumping transitions identified in this work. Microwave spectroscopy in the long-lived $12457.29$-cm-1 state, which we identified as a potential metastable qubit state, demonstrates the feasibility for microwave-based gates and state preparation.